CHM123L section Instructor Significant Figures in Calculations Complete the calculations and enter the answers rounded to the cocrect number of significant figures. In multiple step calculations, track significant figures but do not round until the end. 1. Round to 2 decimal places: 12 6352 Complete these computations the enter the answer rounded to the correct number of significant figures 2 # of significant figures m answer 102794-619 = 3 #of significant figures in answer iyy4 yrtt 4.29 x 103 + 3.99 x 104 # of significant figures in answer- 0.0019 + 0.0352 # of significant figures m answer- - 9 x102 1.73 x103. the calculation showing the cosrect aumber of 3. Complete the Round only the final answer. Enter answer asing a Ene a the last signifoantRend the final antiser. Apply Rwles for Addition and Subtraction bere Reand the final ansver Do not roand yet. 10.0451 9-0.00273 gl 0.1000 g/ml mL 0.1000 g/ml Appl rules for Maultipication and Dasision tofanih the probiem. 3 126g 4,21733 1694
Answer the following questions in complete sentences 4 From your percent difference calculations in Part 2, idennify the most and least acourate piece of glassware. Were these results consistent with expected results? Why or why not? S A student is required to deliver 4.35 mL. of a liquid. Considering significant figures and accuracy, comment on the appropriateness of using the following lab equipment to complete the task. Provide an explanation for cach piece of glassware in your conclusions 10 ml.graduated cylinder 50 mL.graduated cylinder 10 ml. pipet . Buret 3 , l 3229
Answer #1
  1. 12.63
  2. 102.794-61.9 = 40.894. Rounding off, we get 40.9 as the answer. No.of significant digits in the answer = 3.
    4.29 x 10-3 + 3.99 x 10-4 = 4.69 x 10-3. Number of significant digits in the answer = 3.
    0.0019 + 0.0352 = 0.0371. Rounding off, we get; 0.037. Number of significant digits in the answer = 2.
    9 x 10-2 + 1.73 x 10-3 = 9.173 x 10-2. Rounding off, we get; 9 x 10-2. Number of significant digits in the answer = 1
  3. 0.0451-0.002730.04237 0.10000.4237 mL 0.1000
  4. Question not clear
  5. The usage of;
    • 10 mL graduated cylinder
      10 mL cylinders will mostly have a least count of 0.1 mL. Hence can measure the liquid upto around 4.4 mL within error of only 0.05 mL. This is the most appropriate instrument that may be used.
    • 50 mL graduated cylinder
      They will have greater least counts and shall cause errors more than the 10 mL cylinder.
    • Traditional 10 mL pipet cannot be used for measuring any volumes less than 10 mL itself.
    • Chances of air bubbles getting trapped in the Burets are large. Moreover, Burets are meant to measure the flowing liquid, and hence is not a good option.